# Quantum Shortcut to Adiabaticity for State Preparation in a Finite-Sized Jaynes-Cummings Lattice

### v1

@@ Abstract
In noisy quantum systems, achieving high-fidelity state preparation using the adiabatic approach faces a dilemma: either extending the evolution time to reduce diabatic transitions or shortening it to mitigate decoherence effects. Here, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized Jaynes-Cummings lattice by applying a counter-diabatic (CD) driving along given adiabatic trajectories. Leveraging the symmetry of eigenstates in this system, we derive a simplified CD Hamiltonian that only involves local qubit-cavity couplings for a two-site lattice with one polariton excitation. Additionally, we derive the analytical form of the CD Hamiltonian for this lattice with two excitations. Our numerical results demonstrate that this scheme is robust against circuit errors and environmental noise, with characterization achievable through qubit detection. The simplified CD Hamiltonian can be implemented in physical systems with realistic parameters. This approach can lead to a promising pathway to high-fidelity state preparation within a significantly reduced timescale compared to conventional adiabatic methods.
@@ I Introduction
Adiabatic evolution involves a gradual transformation of a quantum system from an initial state to a desired final state by slowly varying a time-dependent Hamiltonian, and has been extensively studied for quantum state preparation in quantum information processing [1, 2]. In noisy quantum systems [3, 4], achieving high fidelity for the prepared states using the adiabatic approach requires a balance in choosing the evolution time: it should be long enough to reduce unwanted diabatic transitions yet short enough to mitigate decoherence from environmental noise. Various approaches have been used to develop quantum shortcut to adiabaticity, where reverse engineering of diabatic transitions is employed to accelerate a slow adiabatic process via nonadiabatic shorcuts [5, 6]. These approaches include counter-diabatic (CD) driving [7, 8, 9, 10, 11], invariant-based approaches [12, 13], the derivative removal of adiabatic gate (DRAG) approach [14], dressed state approaches [15] and etc.. Quantum shortcut to adiabaticity has been proposed and experimentally demonstrated in many quantum systems, such as superconducting qubits, defect qubits, quantum dot arrays, and cold atoms [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. One hurdle that prevents the general application of the quantum shortcut approaches is the complexity of the Hamiltonian required for reverse engineering, which often involves multipartite or nonlocal interactions. Approximate approaches have been explored to overcome this hurdle, including mean-field approximation, variational approach of the effective Hamiltonian, and digitalized approaches [29, 30, 31, 32, 33].
@@
In this work, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized JC lattice by applying CD driving during an adiabatic evolution. In our approach, we explore the symmetry of the system’s eigenstates to simplify the CD Hamiltonians for two-site and three-site lattices. We find that for a two-site lattice with one excitation, the CD Hamiltonian can involve only local qubit-cavity couplings, and with two excitations, the CD Hamiltonian on a subspace of the system can have an analytical form involving four-operator couplings. We also study the effects of control errors and environmental noise on this scheme. Our numerical results demonstrate that this scheme is robust against circuit errors and environmental noise and can be characterized by measuring the qubits. This work hence presents an implementable CD Hamiltonian that shortens the evolution time of an adiabatic process in a JC lattice and provides a promising avenue to achieving high-fidelity state preparation within a significantly reduced timescale.
@@ III Counter-Diabatic Driving
CD driving can be applied to the Hamiltonian $H_{\rm r}(t)$ to cancel diabatic transitions and generate quantum shortcut to adiabaticity during an adiabatic evolution. The CD Hamiltonian has the general form [7, 8, 5]:
@@
In this section, we derive and simplify the CD Hamiltonian for two-site and three-site JC lattices evolving through given adiabatic trajectories. Our results show that the simplified CD Hamiltonian can involve only local qubit-cavity coupling and can be implementable using current technology.
@@ III.1 Two sites one excitation
The eigenstate $v_{1}$ is the ground state with energy $E_{1}$. From (4), we observe that the states $v_{1}$ and $v_{2}$ are symmetric with respect to an exchange of the states on the two sites, while the states $v_{3}$ and $v_{4}$ are antisymmetric and acquire a factor of $-1$ when the states on the two sites are swapped.
@@
Figure 2: Eigenenergies $E_{n}$ of a two-site JC lattice with one excitation. (a) $E_{n}$ vs hopping rate $J$ for $g=1$ and $\Delta=1$. (b) $E_{n}$ vs qubit-cavity coupling $g$ for $J=2$ and $\Delta=1$. All parameters are given in dimensionless units.
@@
In Fig. 2(a), we plot the energy spectrum of the system as a function of the photon hopping rate $J$ for $g=1$ and $\Delta=1$. At $J=0$, $\Delta_{J}^{\pm}=\Delta$, resulting in $E_{1}=E_{3}$ ($E_{2}=E_{4}$), which indicates that the states $v_{1}$ and $v_{3}$ ($v_{2}$ and $v_{4}$) are degenerate states for arbitrary values of the qubit-cavity coupling and detuning. In Fig. 2(b), we plot the energy spectrum vs the qubit-cavity coupling $g$ for $J=2$ and $\Delta=1$. At $g=0$, $E_{2}=E_{3}=0$ when $\Delta<J$.
@@
We want to emphasize that the symmetry of these eigenstates is crucial for our development of a simplified CD Hamiltonian that exclusively involves local qubit-cavity couplings, making it feasible for implementation in practical systems. This symmetry also facilitates the preparation of the ground state $v_{1}$ at $J=0$ via resonant pumping, without inducing excitation to the degenerate state $v_{3}$. Furthermore, during an adiabatic evolution, the variation of the Hamiltonian does not transition the system from a symmetric to an antisymmetric state due to the inherent symmetry of the Hamiltonian.
@@
Ramping of hopping rate $J$. Assuming that the hopping rate $J(t)$ is continuously tuned from $J(0)=0$ at time $t=0$ to $J(T)=J_{f}$ at the final time $T$ with $J_{f}$ being the target hopping rate, while keeping the qubit-cavity coupling and the detuning unchanged during the adiabatic process. At $t=0$, the system is prepared in the ground state $v_{1}$ for $J=0$ by resonantly pumping the system from the state $|0,g\rangle_{1}|0,g\rangle_{2}$. The variation of $J(t)$ induces off-diagonal matrix elements between the instantaneous eigenstates, resulting in undesired transitions to the excited states. The nonzero matrix elements are: $|\langle v_{1}|V_{J}|v_{2}\rangle|=g/\chi_{1}^{-}$, $|\langle v_{3}|V_{J}|v_{4}\rangle|=g/\chi_{1}^{+}$, along with their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$, and between the states $v_{3}$ and $v_{4}$, respectively. We then derive the CD Hamiltonian using (1):
$H_{1}(t)$ $=$ $i\frac{g}{(\chi_{n}^{-})^{2}}\frac{dJ}{dt}(|v_{1}\rangle\langle v_{2}|-|v_{2}\rangle\langle v_{1}|)$ (5)
$-i\frac{g}{(\chi_{n}^{+})^{2}}\frac{dJ}{dt}(|v_{3}\rangle\langle v_{4}|-|v_{4}\rangle\langle v_{3}|).$
@@
Hence, when the system is in the ground state $|v_{1}\rangle$, the only allowable transition is to the symmetric eigenstate $|v_{2}\rangle$. As discussed above, this is due to the symmetry of the eigenstates and the Hamiltonian. A variation of the hopping rate in the Hamiltonian preserves the symmetry of the state, with the only possible diabatic transition being the excitation to the state $|v_{2}\rangle$. Consequently, the state at an arbitrary time $t$ during the evolution remains symmetric and can only be a superposition of the instantaneous $|v_{1}\rangle$ and $|v_{2}\rangle$. Below, we will simplify the CD Hamiltonian using this property.
@@
We first convert the CD Hamiltonian (5) to physical operators of the qubits and cavities in order to discuss its implementation. In the Hilbert space for one polariton excitation, we have $a_{1}^{\dagger}\sigma_{1-}=|1,g\rangle_{1}\langle 0,e|\otimes|0,g\rangle_{2}\langle 0,g|$, $a_{1}^{\dagger}a_{2}=|1,g\rangle_{1}\langle 0,g|\otimes|0,g\rangle_{2}\langle 1,g|$, and so forth. Using (4), we derive:
@@
with $A_{\pm}=a_{1}\pm a_{2}$ and $S_{\pm}=\sigma_{1-}\pm\sigma_{2-}$. The CD Hamiltonian can then be written as
$H_{1}(t)$ $=$ $i\frac{g}{2(\chi_{n}^{-})^{2}}\frac{dJ}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (7)
$-i\frac{g}{2(\chi_{n}^{+})^{2}}\frac{dJ}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
@@
This Hamiltonian comprises local couplings between qubits and cavities on the same site, such as $\sigma_{1+}a_{1}$ and $\sigma_{2+}a_{2}$, as well as nonlocal couplings between qubits and neighboring cavities, such as $\sigma_{1+}a_{2}$ and $\sigma_{2+}a_{1}$.
As seen from (5), when the initial state is the ground state $v_{1}$, the CD Hamiltonian only eliminates excitation from $v_{1}$ to $v_{2}$. Thus, the second term in $H_{1}(t)$ does not actively contribute to the elimination of diabatic transitions, and its amplitude can be adjusted at will without affecting the adiabatic evolution. We then change the coefficient of the second term in (7) from $-ig/2(\chi_{n}^{+})^{2}$ to $ig/2(\chi_{n}^{-})^{2}$, and obtain the simplified CD Hamiltonian
$H_{1}^{\prime}(t)=ig^{\prime}\left(a_{1}\sigma_{1}^{+}-a_{1}^{\dagger}\sigma_{1}^{-}+a_{2}\sigma_{2}^{+}-a_{2}^{\dagger}\sigma_{2}^{-}\right).$ (8)
@@
The Hamiltonian $H_{1}^{\prime}$ includes only local qubit-cavity couplings with a strength given by $g^{\prime}=\frac{g}{(\chi_{n}^{-})^{2}}\frac{dJ}{dt}$, which can be interpreted as a Jaynes-Cummings coupling with a $e^{i\pi/2}$ phase. This Hamiltonian can be engineered by adjusting the local qubit-cavity couplings.
@@
For demonstration, we simulate the time evolution of this system under the adiabatic Hamiltonian $H_{\rm r}$ with either the CD Hamiltonian $H_{1}$ or $H_{1}^{\prime}$ applied. We investigate both linear ramping of the hopping rate, with $J(t)=J_{f}(t/T)$, and quadratic ramping, with $J(t)=J_{f}(t/T)^{2}$. For linear ramping, $dJ/dt=J_{f}/T$; and for quadratic ramping, $dJ/dt=2J_{f}t/T^{2}$. We choose the target hopping rate $J_{f}=2$, the qubit-cavity coupling $g\equiv 1$, the detuning $\Delta\equiv 1$, and the total time $T=0.5\pi$. We define the fidelity of the quantum state at time $t$ as $F(t)=|\langle v_{1}(t)|v(t)\rangle|^{2}$, where $|v(t)\rangle$ is the state at time $t$ and $|v_{1}(t)\rangle$ is the instantaneous ground state of the adiabatic Hamiltonian $H_{\rm r}$ at time $t$. In Fig. 3(a), we plot the infidelity $1-F(t)$ vs $t/T$. Without the CD Hamiltonian, the infidelity increases significantly during the evolution, as $T=0.5\pi$ is a very short time and diabatic transitions can be serious. In contrast, when the CD Hamiltonian is applied, the infidelity remains below $10^{-10}$ throughout the evolution, preserving the system in the ground state. Both CD Hamiltonians $H_{1}$ and $H_{1}^{\prime}$ yield identical results, confirming our analysis. In Fig. 3(b), we plot the infidelity $1-F(T)$ of the final state vs the total evolution time $T$. For short evolution times, the infidelity of the adiabatic process is quite high. However, as $T$ increases, the infidelity decreases to below $10^{-3}$. These numerical results demonstrate that the CD approach can ensure high fidelity for short evolution times, which is crucial for quantum devices in the NISQ regime.
@@
Figure 3: (Color online) (a) Infidelity $1-F(t)$ vs $t/T$ for a total evolution time of $T=0.5\pi$. (b) Infidelity $1-F(T)$ vs $T$. All solid lines: adiabatic evolution under the Hamiltonian $H_{\rm r}$; and all circles: adiabatic evolution with the CD Hamiltonian $H_{1}^{\prime}$ applied. Blue (red): linear (quadratic) ramping of the hopping rate $J$ for $g\equiv 1$, $\Delta\equiv 1$, $J(0)=0$ and $J_{f}=2$; and green: linear ramping of the qubit-cavity coupling $g$ for $J\equiv 2$, $\Delta\equiv 1$, $g(0)=0$ and $g_{f}=1$.
@@
Ramping of qubit-cavity coupling $g$. Now consider an adiabatic trajectory where the qubit-cavity coupling is ramped linearly with $g(t)=g_{f}t/T$ and $g_{f}$ being the target coupling strength. The nonzero transition matrix elements caused by the variation of $g$ are $|\langle v_{1}|V_{g}|v_{2}\rangle|=\Delta_{J}^{-}/\chi_{1}^{-}$, $|\langle v_{3}|V_{g}|v_{4}\rangle|=\Delta_{J}^{+}/\chi_{1}^{+}$, and their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$, as well as between the states $v_{3}$ and $v_{4}$, due to the symmetry of the eigenstates. The CD Hamiltonian can be derived as:
$H_{1}(t)$ $=$ $i\frac{\Delta_{J}^{-}}{2(\chi_{n}^{-})^{2}}\frac{dg}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (9)
$-i\frac{\Delta_{J}^{+}}{2(\chi_{n}^{+})^{2}}\frac{dg}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
@@
Similar to (7), this Hamiltonian includes couplings between qubits and their neighboring cavities. As discussed above, when $|v_{1}\rangle$ is the initial state, only the state $|v_{2}\rangle$ could be excited by diabatic transitions. Therefore, we can vary the coefficient of the second term in (9). By changing this coefficient from $-i\Delta_{J}^{+}/2(\chi_{n}^{+})^{2}$ to $i\Delta_{J}^{-}/2(\chi_{n}^{-})^{2}$, we obtain a simplified CD Hamiltonian:
@@
which only involves local qubit-cavity couplings with strength $g^{\prime}=\frac{\Delta_{J}^{-}}{(\chi_{1}^{-})^{2}}\frac{dg}{dt}$. Consequently, implementing both the adiabatic Hamiltonian and the CD Hamiltonian in this trajectory only requires tuning the local qubit-cavity couplings. This feature can significantly simplify the experimental setup.
We perform a numerical simulation of this process, choosing a hopping rate of $J\equiv 2$, a detuning of $\Delta\equiv 1$, and a target coupling strength of $g_{f}=1$. At $t=0$, with $g(0)=0$, the initial state is $(|1,g\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,g\rangle_{2})/\sqrt{2}$, where the excitation is stored in the cavity modes. The infidelity $1-F(t)$ vs $t/T$ for $T=0.5\pi$ is plotted in Fig.3(a), and the infidelity at the final time $1-F(T)$ vs the total evolution time $T$ is plotted in Fig.3(b). The results exhibit similar behavior to the ramping of the hopping rate $J$.
Self-protected trajectory. The magnitude of the CD Hamiltonian in (10) is proportional to $\Delta_{J}^{-}=\Delta-J$. In the special case of $\Delta=J$, $\Delta_{J}^{-}=0$, leading to the interesting result of $H_{1}^{\prime}=0$. This indicates no CD Hamiltonian is needed to eliminate the diabatic transitions. It can be shown that the matrix element $|\langle v_{1}|V_{g}|v_{2}\rangle|\equiv 0$ under this condition, thus, the only allowable diabatic transition from the ground state to the excited states disappears. The Hamiltonian $H_{\rm r}$ is hence self-protected from diabatic transitions, meaning there is no diabatic transition to the excited states caused by ramping the system parameters. The ground state remains in the state $|v_{1}\rangle=\frac{1}{\sqrt{2}}\left(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2}\right)$ throughout the entire evolution, with the excitation occupying the lower polariton state $|1,-\rangle$ in one of the two lattice sites. This is confirmed by our numerical simulation of this trajectory.

### v2

@@ Abstract
In noisy quantum systems, achieving high-fidelity state preparation using the adiabatic approach faces a dilemma: either extending the evolution time to reduce diabatic transitions or shortening it to mitigate decoherence effects. Here, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized Jaynes-Cummings lattice by applying counter-diabatic (CD) driving along given adiabatic trajectories. Leveraging the symmetry of eigenstates in our system, we convert the CD driving to an implementable Hamiltonian that only involves local qubit-cavity couplings for a two-site lattice with one polariton excitation. Additionally, we derive a partial analytical form of the CD driving for the lattice with two excitations. Our numerical results demonstrate that circuit errors and environmental noise have negligible effects on our scheme under practical parameters. We also show that our scheme can be characterized through the detection of qubit operators. This approach can lead to a promising pathway to high-fidelity state preparation in a significantly reduced timescale when compared to conventional adiabatic methods.
@@ I Introduction
Adiabatic evolution involves a gradual transformation of a quantum system from an initial state to a desired final state by slowly varying a time-dependent Hamiltonian, and has been extensively studied for quantum state preparation in quantum information processing [1, 2]. In noisy quantum systems [3, 4], achieving high fidelity for the prepared states using the adiabatic approach requires a balance in choosing the evolution time: it should be long enough to reduce unwanted diabatic transitions yet short enough to mitigate decoherence from environmental noise. Various approaches have been used to develop quantum shortcut to adiabaticity, where reverse engineering of diabatic transitions is employed to accelerate a slow adiabatic process via nonadiabatic shorcuts [5, 6, 7, 8]. These approaches include counter-diabatic (CD) driving [9, 10, 11, 12, 13], invariant-based approaches [14, 15], the derivative removal of adiabatic gate (DRAG) approach [16], dressed state approaches [17] etc.. Quantum shortcut to adiabaticity has been proposed and experimentally demonstrated in many quantum systems, such as superconducting qubits, defect qubits, quantum dot arrays, and cold atoms [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. One hurdle that prevents the general application of these quantum shortcut approaches is the complexity of the Hamiltonian required for reverse engineering, which often involves multipartite or nonlocal interactions. A number of approaches have been explored to overcome this hurdle by deriving local CD driving, including approach that explores the self-similar nature of a quantum system, mean-field approximation, variational approach, Floquet engineering, digitalized approach, and optimal control approach [31, 32, 33, 34, 38, 35, 36, 37, 39, 40].
@@
In our current work, we present a quantum shortcut to adiabaticity for state preparation in a finite-sized JC lattice by applying CD driving during an adiabatic evolution. In our approach, we explore the symmetry of the system’s eigenstates to convert the CD driving to an implementable Hamiltonian for two-site and three-site lattices. We find that for a two-site lattice with one excitation, the CD driving can involve only local qubit-cavity couplings, and with two excitations, the CD driving can have a partial analytical form that involves four-operator couplings in the antisymmetric subspace. We also study the effects of control errors and environmental noise on our scheme, which complements previous CD driving studies on noisy and open quantum systems [36, 64, 65]. Our numerical results demonstrate that the effects of circuit errors and environmental noise are negligible under practical parameters. Meanwhile, this scheme can be characterized by performing measurement on the qubits. This work hence presents an implementable CD driving that can greatly shorten the evolution time of an adiabatic process in a JC lattice and provides a promising avenue to achieving high-fidelity state preparation in a significantly reduced timescale.
@@ III Counter-Diabatic Driving
CD driving can be applied to the Hamiltonian $H_{\rm r}(t)$ to cancel diabatic transitions and generate quantum shortcut to adiabaticity during an adiabatic evolution. The CD driving has the general form [9, 10, 5]:
@@
In this section, we convert the CD driving in (1) to an implementable Hamiltonian for adiabatic evolutions in two-site and three-site JC lattices. Our results show that the CD driving can comprise only local qubit-cavity couplings and is implementable using current technology.
@@ III.1 Two sites one excitation
The eigenstate $v_{1}$ is the ground state with energy $E_{1}$. From (4), we observe that the states $v_{1}$ and $v_{2}$ are symmetric with respect to the exchange of the states on the two sites, while the states $v_{3}$ and $v_{4}$ are antisymmetric and acquire a factor of $-1$ when the states on the two sites are swapped.
@@
Figure 2: Eigenenergies $E_{n}$ ($n\in[1,4]$) of a two-site JC lattice with one excitation. (a) $E_{n}$ vs hopping rate $J$ for $g=1$ and $\Delta=1$. (b) $E_{n}$ vs coupling constant $g$ for $J=2$ and $\Delta=1$. All parameters are in dimensionless units.
@@
In Fig. 2(a), we plot the energy spectrum of the system as a function of the photon hopping rate $J$ for $g=1$ and $\Delta=1$. At $J=0$, $\Delta_{J}^{\pm}=\Delta$, resulting in $E_{1}=E_{3}$ ($E_{2}=E_{4}$), which indicates that the states $v_{1}$ and $v_{3}$ ($v_{2}$ and $v_{4}$) are degenerate. In Fig. 2(b), we plot the energy spectrum vs the qubit-cavity coupling $g$ for $J=2$ and $\Delta=1$. At $g=0$ and $\Delta<J$, $E_{2}=E_{3}=0$, and $v_{2}$ and $v_{3}$ are degenerate.
@@
We want to emphasize that the symmetry of these eigenstates is crucial for our development of an implementable CD driving that only involves local qubit-cavity couplings. During an adiabatic evolution, the variation of the Hamiltonian does not induce transition of the system from a symmetric to an antisymmetric state due to the inherent symmetry of the Hamiltonian. Furthermore, this symmetry also facilitates the preparation of the ground state $v_{1}$ at $J=0$ via resonant pumping, without inducing excitation to the state $v_{3}$ that is degenerate to $v_{1}$.
@@
Ramping of hopping rate $J$. Assuming that the hopping rate $J(t)$ is continuously tuned from $J(0)=0$ at time $t=0$ to $J(T)=J_{f}$ at the final time $T$ with $J_{f}$ being the target hopping rate, while keeping the qubit-cavity coupling and the detuning unchanged during the adiabatic process. At $t=0$, the system is prepared in the ground state $v_{1}$ for $J=0$ by resonantly pumping the system from the state $|0,g\rangle_{1}|0,g\rangle_{2}$. The variation of $J(t)$ induces off-diagonal matrix elements between the instantaneous eigenstates, resulting in undesired transitions to the excited states. The nonzero matrix elements are: $|\langle v_{1}|V_{J}|v_{2}\rangle|=g/\chi_{1}^{-}$, $|\langle v_{3}|V_{J}|v_{4}\rangle|=g/\chi_{1}^{+}$, along with their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$, and between the states $v_{3}$ and $v_{4}$, respectively. We then derive the CD driving using (1):
$H_{1}(t)$ $=$ $i\frac{g}{(\chi_{1}^{-})^{2}}\frac{dJ}{dt}(|v_{1}\rangle\langle v_{2}|-|v_{2}\rangle\langle v_{1}|)$ (5)
$-i\frac{g}{(\chi_{1}^{+})^{2}}\frac{dJ}{dt}(|v_{3}\rangle\langle v_{4}|-|v_{4}\rangle\langle v_{3}|).$
@@
Hence, when the system is in the ground state $v_{1}$, the only allowable transition is to the symmetric eigenstate $v_{2}$. As discussed above, this is due to the symmetry of the eigenstates and the Hamiltonian. A variation of the hopping rate in the Hamiltonian preserves the exchange symmetry of the state, which can only induce diabatic transition to the state $v_{2}$. Consequently, the state at an arbitrary time $t$ during the evolution remains symmetric and can only be a superposition of the instantaneous $v_{1}$ and $v_{2}$ states. Thus, the second term in $H_{1}(t)$ does not actively contribute to the elimination of diabatic transitions, and its amplitude can be adjusted with flexibility, without affecting the adiabatic dynamics of this system. Below, we will construct an easier-to-implement CD driving using this property.
@@
We first convert the CD driving in (5) to physical operators of the qubits and cavities to see if it can be easily implemented. In the Hilbert space for one polariton excitation, we have $a_{1}^{\dagger}\sigma_{1-}=|1,g\rangle_{1}\langle 0,e|\otimes|0,g\rangle_{2}\langle 0,g|$, $a_{1}^{\dagger}a_{2}=|1,g\rangle_{1}\langle 0,g|\otimes|0,g\rangle_{2}\langle 1,g|$, and so forth. Using the expression of the eigenstates in (4), we derive:
@@
with $A_{\pm}=a_{1}\pm a_{2}$ and $S_{\pm}=\sigma_{1-}\pm\sigma_{2-}$. And the CD driving is
$H_{1}(t)$ $=$ $i\frac{g}{2(\chi_{1}^{-})^{2}}\frac{dJ}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (7)
$-i\frac{g}{2(\chi_{1}^{+})^{2}}\frac{dJ}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
@@
This Hamiltonian comprises local couplings between qubits and cavities on the same site such as $\sigma_{1+}a_{1}$ and $\sigma_{2+}a_{2}$ as well as nonlocal couplings between qubits and neighboring cavities such as $\sigma_{1+}a_{2}$ and $\sigma_{2+}a_{1}$.
By changing the coefficient of the second term in (7) from $-ig/2(\chi_{1}^{+})^{2}$ to $ig/2(\chi_{1}^{-})^{2}$, we obtain the following CD Hamiltonian:
$H_{1}^{\prime}(t)=ig^{\prime}\left(a_{1}\sigma_{1+}-a_{1}^{\dagger}\sigma_{1-}+a_{2}\sigma_{2+}-a_{2}^{\dagger}\sigma_{2-}\right).$ (8)
@@
This Hamiltonian includes only local qubit-cavity couplings with coupling strength given by $g^{\prime}=\frac{g}{(\chi_{1}^{-})^{2}}\frac{dJ}{dt}$. It can be interpreted as a Jaynes-Cummings coupling with a $e^{i\pi/2}$ phase, and can be implemented by adjusting local qubit-cavity couplings.
@@
For demonstration, we simulate the time evolution of this system under the adiabatic Hamiltonian $H_{\rm r}$ with either $H_{1}$ or $H_{1}^{\prime}$ applied. We consider both linear ramping of the hopping rate with $J(t)=J_{f}\frac{t}{T}$ and quadratic ramping of the hopping rate with $J(t)=J_{f}\left(\frac{t}{T}\right)^{2}$. For linear ramping, $dJ/dt=J_{f}/T$; and for quadratic ramping, $dJ/dt=2J_{f}t/T^{2}$. We choose the target hopping rate $J_{f}=2$, the qubit-cavity coupling $g\equiv 1$, the detuning $\Delta\equiv 1$, and the total time $T=0.5\pi$. We define the fidelity of the quantum state at time $t$ as $F(t)=|\langle v_{1}(t)|v(t)\rangle|^{2}$, where $v(t)$ is the state at time $t$ and $v_{1}(t)$ is the instantaneous ground state of the adiabatic Hamiltonian at time $t$. In Fig. 3(a), we plot the infidelity $1-F(t)$ vs $t/T$. Without the CD driving, the infidelity increases significantly during the evolution, as the diabatic transitions can be serious for a very short evolution time of $T=0.5\pi$. In contrast, when the CD driving $H_{1}^{\prime}$ is applied, the infidelity, up to a small numerical error below $10^{-10}$, is zero throughout the evolution, and the system is preserved in the ground state. In Fig. 3(b), we plot the infidelity $1-F(T)$ of the final state vs the total evolution time $T$. Without the CD driving, the infidelity can be quite high for a short evolution time. This result demonstrates that the CD approach can ensure high fidelity for a short evolution time, which is crucial for devices in the noisy intermediate-scale quantum (NISQ) regime [3]. As shown in Fig. 7(a, b) in Appendix A, the infidelities under $H_{1}$ and $H_{1}^{\prime}$ are the same up to a small numerical error below $10^{-10}$, which confirms our analysis that these two CD drivings yield the same dynamics for the initial state $v_{1}$.
@@
Figure 3: (Color online) (a, c) Infidelity $1-F(t)$ vs $t/T$ for a total evolution time of $T=0.5\pi$ for two-site and three-site JC lattices, respectively. (b, d) Infidelity $1-F(T)$ vs $T$ for two-site and three-site JC lattices, respectively. In all plots, blue solid line (blue circles): linear ramping of hopping rate $J$ for $g\equiv 1$, $\Delta\equiv 1$, $J(0)=0$ and $J_{f}=2$ under the adiabatic Hamiltonian $H_{\rm r}$ only (with the CD driving $H_{1}^{\prime}$ applied); red dashed line (red squares): quadratic ramping of $J$ for $g\equiv 1$, $\Delta\equiv 1$, $J(0)=0$ and $J_{f}=2$ under $H_{\rm r}$ (with $H_{1}^{\prime}$); and green dotted line (green triangles): linear ramping of coupling $g$ for $J\equiv 2$, $\Delta\equiv 1$, $g(0)=0$ and $g_{f}=1$ under $H_{\rm r}$ (with $H_{1}^{\prime}$). All parameters are in dimensionless units.
@@
Ramping of qubit-cavity coupling $g$. Now consider an adiabatic trajectory where the qubit-cavity coupling is ramped linearly with $g(t)=g_{f}t/T$ and $g_{f}$ being the target coupling strength. The nonzero transition matrix elements caused by the variation of $g$ are $|\langle v_{1}|V_{g}|v_{2}\rangle|=\Delta_{J}^{-}/\chi_{1}^{-}$, $|\langle v_{3}|V_{g}|v_{4}\rangle|=\Delta_{J}^{+}/\chi_{1}^{+}$, and their conjugate elements, which induce transitions between the states $v_{1}$ and $v_{2}$ as well as between the states $v_{3}$ and $v_{4}$, respectively, due to the symmetry of the eigenstates. The CD driving can be derived as:
$H_{1}(t)$ $=$ $i\frac{\Delta_{J}^{-}}{2(\chi_{1}^{-})^{2}}\frac{dg}{dt}\left(S_{+}^{\dagger}A_{+}-A_{+}^{\dagger}S_{+}\right)$ (9)
$-i\frac{\Delta_{J}^{+}}{2(\chi_{1}^{+})^{2}}\frac{dg}{dt}\left(S_{-}^{\dagger}A_{-}-A_{-}^{\dagger}S_{-}\right).$
@@
Similar to (7), this Hamiltonian includes nonlocal couplings between qubits and neighboring cavities. As discussed above, when $|v_{1}\rangle$ is the initial state, only the state $|v_{2}\rangle$ could be excited by diabatic transitions. Therefore, we can vary the coefficient of the second term in (9). By changing this coefficient from $-i\Delta_{J}^{+}/2(\chi_{1}^{+})^{2}$ to $i\Delta_{J}^{-}/2(\chi_{1}^{-})^{2}$, we obtain the CD driving:
@@
which only involves local qubit-cavity couplings with strength $g^{\prime}=\frac{\Delta_{J}^{-}}{(\chi_{1}^{-})^{2}}\frac{dg}{dt}$. Consequently, implementing both the adiabatic Hamiltonian and the CD driving only requires the tuning of the local qubit-cavity coupling, which can significantly simplify the experimental setup.
We perform a numerical simulation on this system with hopping rate of $J\equiv 2$, detuning of $\Delta\equiv 1$, and a target coupling strength of $g_{f}=1$. At $t=0$ and $g(0)=0$, the initial state is $(|1,g\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,g\rangle_{2})/\sqrt{2}$, where the excitation is stored in the cavity modes. The infidelity $1-F(t)$ vs $t/T$ for $T=0.5\pi$ is plotted in Fig.3(a), and the infidelity at the final time $1-F(T)$ vs the total time $T$ is plotted in Fig.3(b). The results are similar to the case of ramping the hopping rate $J$.
Self-protected trajectory. The magnitude of the CD driving in (10) is proportional to $\Delta_{J}^{-}=\Delta-J$. In the special case of $\Delta=J$, $\Delta_{J}^{-}=0$, leading to the interesting result: $H_{1}^{\prime}=0$, which means that no CD driving is required. It can be shown that the matrix element $|\langle v_{1}|V_{g}|v_{2}\rangle|\equiv 0$ under this condition, hence the only allowable diabatic transition for the initial state $v_{1}$ disappears. The Hamiltonian $H_{\rm r}$ is thus self-protected from diabatic transition, i.e., there is no diabatic transition when ramping the qubit-cavity coupling $g$. The ground state is $\frac{1}{\sqrt{2}}\left(|1,-\rangle_{1}|0,g\rangle_{2}+|0,g\rangle_{1}|1,-\rangle_{2}\right)$ throughout the evolution, with the system occupying the lower polariton state $|1,-\rangle$ in one of the lattice sites.